### v2

# Good Old-Fashioned Artificial Intelligence meets Generative Artificial Intelligence: Development of Expert Systems by means of Large Language Models

Abstract
The development of large language models (LLMs) has successfully transformed knowledge-based systems such as open domain question nswering, which can automatically produce vast amounts of seemingly coherent information. Yet, those models have several disadvantages like hallucinations or confident generation of incorrect or unverifiable facts. In this paper, we introduce a new approach to the development of expert systems using LLMs in a controlled and transparent way. By limiting the domain and employing a well-structured prompt-based extraction approach, we produce a symbolic representation of knowledge in Prolog, which can be validated and corrected by human experts. This approach also guarantees interpretability, scalability and reliability of the developed expert systems. Via quantitative and qualitative experiments with Claude Sonnet 3.7 and GPT-4.1, we show strong adherence to facts and semantic coherence on our generated knowledge bases. We present a transparent hybrid solution that combines the recall capacity of LLMs with the precision of symbolic systems, thereby laying the foundation for dependable AI applications in sensitive domains.
1 Introduction
Since the firsts expert systems were developed in the 70’s and 80’s like DENDRAL [8, 10], MYCIN [30], XCON [2], and some others, this technique has not stopped evolving as an aid to knowledge and advice in a multitude of tasks like medical, legal, teaching, etc…
One of the greatest complexities when creating an expert system is to provide it with complete and reliable knowledge. For this task, a human expert knowledgeable in the subject and numerous questionnaires or other methods were generally used to capture the knowledge obtained in graphs, decision trees, knowledge rules [13] in order to squeeze the knowledge obtained in the best possible way [23].
With the appearance in 2010 of LLM’s (large language models), systems with a huge information capacity of data, a whole world of possibilities opens when it comes to generating systems with a smaller specific domain aimed for a particular task. Thus, chatbots for medicine, education, and completion have been created in no time, allowing fluid dialogs and multitude of new questions to be answered [9]. The problem is that not all the knowledge returned by the system is reliable. As indicated by [15] and [18] many times the LLM in its eagerness to give an answer produces hallucinations, i.e. wrong or not accurate knowledge as an answer. The problem here is serious, because if the sources of knowledge are not reliable, the answers should not be reliable either, so, unless we check it, the system will be spreading wrong information.
Hallucinations within an LLM are false responses that the system produces when there is either inaccurate data or an internal failure. Detecting these hallucinations is critical when implementing an LLM in a system, since it can induce serious errors in sensitive processes, such as medical, scientific, business, etc., and can contribute to the dissemination of false information in the networks due to lack of verification or interpretation [24].
The generation of hallucinations can be categorized as a failure for any of these reasons [27, 20]: Data failure [7], because they are not up to date, there is biased data or they are mixed with fake news, intellectual property problems when accessing certain information sources, etc. In the case proposed in Figure LABEL:fig:papa, there is a clear case of outdated information, since the system did not have in its data the death of Pope Francis 15 days before the question was asked. Even so, it asserts that as of May 7, the pope is still alive. To solve this problem, it is advisable to use reliable sources of information, contrasted information (for example in the case of medical issues, the Mayo Clinic has extensive information on its website https://www.mayoclinic.org/es), and if the information is not accessible or there is not enough information to formulate an adequate answer, do not answer the question.
Failure in the model and training, architecture problems, RNN’s vs Transformers, problems in understanding the context and complexity of the question, overfitting problems in the data training to generate an answer.
And finally, problems in the generation of the answer because many times LLMs focus more on the fluency of the presented text than on the accuracy of the information. A high sampling temperature increases randomness, while focusing on the words with the highest probability can result in a loss of text quality [26].
There are several proposals to solve these problems, including using reliable external sources, doing fact-checking within the application itself to check that it shows the same results (using synonyms, different temperatures, etc.) or training the systems to detect erroneous answers [14, 25].
To propose another solution to this problem, we have designed a system that addresses a smaller and more specific knowledge domain. Through a of a set of predefined prompts [16, 1, 19, 12] a LLM such as Chatgpt can generate a more manageable knowledge base to be transformed into an easily verifiable model (such as a system of rules [32]), by a human expert in the field. By means of this verification, we guarantee three very important advantages when generating the expert system:
  • Explainability of the system.
  • Greater volume of information (since it is easier to produce information in this way than through an expert).
  • Veracity and reliability of the information (since if the expert detects any error in the model presented, it will be corrected immediately).
Regarding the creation of an expert system, Prolog is a well-known logic programming language based on first-order predicate logic and has long remained popular for the development of expert systems due to its declarative pattern matching syntax and inference mechanism. System knowledge is expressed by facts and rules, and the system deduces new knowledge using backward chaining and resolution. Famous applications are medical diagnosis systems (e.g., MYCIN), legal reasoning tools [3], and configuration of complex products [28, 5]. For example, Prolog was the programming language employed by the clinical question answering system CLINIQA [22] and by XCON (also called R1), a rule-based system used at Digital Equipment Corporation to specify how to configure machines [21]. Its capacity of dealing with symbolic reasoning makes it an appropriate tool for constructing tutoring systems and intelligent advisors [11, 6]. Prolog is still a very good option to program expert systems in the fields which require explainability, transparency and rule based reasoning.
To achieve these goals, we have built a model that queries the information from the LLM though a set of specific prompts (designed by an expert), then generates a model of logical rules to ease the expert the validation phase, a correction method to clean the information from imprecisions, wrong data or data bias, and the creation of a reliable expert system. So, the paper is divided as follows. First, we start by introducing how expert systems have evolved since the 1970s and the role that Large Language Models (LLMs) have played in a new era of knowledge extraction for systems such as chatbots and recommendation engines. We have remarked that LLMs tend to hallucinate, generate incorrect or unverifiable outputs, which can cause harmful decisions in important tasks, such as medicine, education, law.
Section preliminary and related works, presents a state of the art of the works that have served as basis for this project, reviewing its achievements and main contributions. In the methodology section, we have developed a protocol to select, define and process the knowledge of an LLM. For this task, we have used customized prompts to extract structured information from language models. This information is represented in Prolog, which excels in transparency, explainability and rule-based reasoning. In the methodology section we have presented a recursive procedure which constructs conceptual graphs interrogating the LLM, translating its response to Prolog facts and relations, and validating the knowledge syntactically and semantically. The experimental section shows both quantitative and qualitative results. We also demonstrate that the Prolog expert systems generated can be successfully executed and queried without errors, confirming their feasibility for practical use. Finally, in the last section, conclusions and future works are presented.
2 Preliminary and related work
In this section, we will review the work that has been developed in this area and has served as the basis for our system. We will divide them into three main parts. On the one hand, information gathering, where we will address the method used to query the LLM information. Secondly, we will focus on the formal representation of the knowledge and its validation, and finally we will address the construction of the expert system.
2.1 Knowledge adquisition
Since 2010, with the appearance of LLMs, LlaMA, Gpt3, Copilot, applications that rely on these systems have not ceased to appear, from recommendation systems, chatbots, classifiers, automatic programmers, etc., have revolutionized the digital world.
Knowledge generation has gone through an algorithmic treatment different from traditional methods, such as fine-tuning [19] or knowledge extraction from the LLM itself through prompts [17]). This method is usually used to reduce the knowledge domain of the LLM itself into a more specific, controllable and transparent one.
For this task, and based on these works, we have designed a complete set of prompts that cover the knowledge domain that we want to test in these experiments.
2.2 Limitations of LLMs regarding verification systems and computability
Even though large language models (LLMs) possess incredible abilities in solving everything from natural language generation to code completion, performance starkly demarcates boundaries put in place by classical computability theory. For example, when prompted to check if an arbitrary piece of code will terminate, LLMs will sometimes provide seemingly convincing rationale or even provide an answer authoritatively. This is not, however, a result of any decision-making property but rather statistical inference from training data. The Halting Problem, which Turing proved undecidable, states that there exists no algorithm—not even implicitly encoded ones in LLMs—capable of deciding termination for all possible programs [4]). The confident prediction by the model is therefore a simulation of understanding rather than actual verification. Analogously, when prompted to prove something in formal logic or do stepwise algebraic manipulations, LLMs often hallucinate intermediate steps or misapply logical principles, demonstrating their lack of formal basis.
It is this epistemic gap which has special resonance in areas where correctness is not a matter of plausibility, but proof, e.g. formal methods, verification of cryptographic protocols, theorem proving. The difficulty is not one of scale but principle: LLMs, bounded as they are by the ChurchTuring thesis, work wholly within the universe of computable functions and therefore share the undecidability and incompleteness theorems which constrain all formal systems of adequately high complexity. What LLMs provide accordingly is a probabilistic shadow of mathematical argument good for heuristic exploration, but incapable in principle of delivering guarantees. In this manner, then, LLMs’ illusion of omniscience may well obscure the very theoretical boundaries that are as relevant in the age of AI as in Gödel’s and Turing’s era.
3 Methodology
This section presents the methodological framework for extracting structured, symbolic knowledge from a large language model (LLM) and encoding it into a Prolog-based expert system. The objective is to construct a formal, logically consistent knowledge base representing the semantic landscape surrounding a user-defined concept. The proposed system combines recursive prompt chaining, ontology-aligned relation extraction, and symbolic encoding in first-order logic.
The pipeline is implemented in Python and invoked via a shell interface that receives a root keyword $T\in K$, where $K$ denotes the set of all conceptual domains. The knowledge acquisition process is driven by two hyper parameters: the horizontal breadth $h\in N$, which limits the number of semantically related concepts retrieved per node, and the vertical depth $d\in N$, which defines the maximum depth of the conceptual graph rooted at $T$. The system interacts with the LLM via structured prompts, receiving a JSON-formatted response with concept nodes and labeled semantic relations, which are translated into Prolog facts using a controlled predicate vocabulary.
Formally, the output is a directed labeled graph $G=(C,R)$, where:
  • $C\subseteq K$ is the set of extracted concepts.
  • $R\subseteq C\times L\times C$ is the set of relations among concepts.
  • $L$ is a finite set of predicate labels, including generic forms (related_to, implies, causes, relates_to) and domain-specific extensions.
To manage graph expansion, we define a recursive semantic expansion operator: $\Psi:C\longrightarrow\mathcal{P}(C\times L\times C)$ which maps a concept to a set of at most $h$ labelled relations, and is recursively applied along newly discovered nodes up to depth $d$. The traversal is breadth-first and avoids cycles via memorization of visited concepts. We summarize these steps in Algorithm 1.
Input: $T$: String // Root concept
    $h$: Integer $\geq 1$ // Horizontal breadth
    $d$: Integer $\geq 0$ // Vertical depth
Output: $K_{T}$: Prolog file // Prolog knowledge base
1:procedure Build_Knowledge_Base($T,h,d$)
2:  $Q\leftarrow[(T,0)]$ $\triangleright$ Queue of (concept, depth)
3:  $V\leftarrow\emptyset$ $\triangleright$ Visited concepts
4:  $F\leftarrow\emptyset$ $\triangleright$ Set of unique fact hashes
5:  $KB\leftarrow\emptyset$ $\triangleright$ Prolog fact accumulator
6: ;
While $Q\neq\emptyset$
7:    $(c,\text{depth})\leftarrow Q.\text{dequeue}()$
8:    If $c\in V$ or $\text{depth}>d$
9:      continue
10:    Endif
11:    $V\leftarrow V\cup\{c\}$
12:    $\text{Json}\leftarrow\text{Query\_LLM}(c,h)$
13:    $(\text{Concepts},\text{Relations})\leftarrow\text{Parse\_JSON}(\text{Json})$
14:    ForAll $c^{\prime}\in\text{Concepts}$
15:      If $c^{\prime}\notin V$
16:        $Q.\text{enqueue}((c^{\prime},\text{depth}+1))$
17:      EndIf
18:      $KB\leftarrow KB\cup\{\text{concept}(c^{\prime}),\text{related\_to}(c^{\prime},c)\}$
19:    EndFor
20:    ForAll $(s,r,t,\text{explanation})\in\text{Relations}$
21:      $KB\leftarrow KB\cup\{r(s,t)\}$
[…]
23:    EndFor
24:    Deduplicate(KB,F)(KB,F)
25: ;
EndWhile
[…]
28:  return $K_{T}$
29:end procedure
Algorithm 1 Recursive Semantic Expansion and Symbolic Encoding
Each extracted concept $c$ is encoded as a unary predicate concept($c$)., and each semantic relation $(c1,r,c2)$ as a binary predicate $r(c1,c2)$., where $r\in L$. Natural language explanations generated by the LLM are preserved as inline Prolog comments using $\%$ Explanation: directives. The system enforces uniqueness through lexical normalization and hashing of facts, and performs syntax validation using SWI-Prolog before deployment.
This algorithm supports the construction of multi-layered conceptual graphs rooted in an arbitrary input topic, enabling recursive reasoning and logical querying within an interpretable and extensible expert system framework.
The system supports both generic and domain-specialized relation extraction, enabling the construction of knowledge graphs that vary in granularity and scope. In its default configuration, the model adapts prompts based on the classified domain of the input topic—such as philosophy, history, or literature—yielding relations like developed_by, lived_in, wrote, pioneered, or invented, which capture biographical, geographical, and authorship information. This allows for richer, discipline-specific knowledge representations. However, the system can be configured to operate in a minimal mode, restricting the ontology to the core predicates concept/1, related_to/2, implies/2, and causes/2, thereby focusing purely on abstract conceptual structure and logical causality. This modularity enables flexible deployment for applications requiring either high interpretability or lightweight logical inference. We describe all the predicates in Table 1 and Table 2.
Table 1: Core Predicates Supported by the System.
Category Predicate Arity Description
Core (Causal) concept/1 1 Declares a concept
related_to/2 2 Links a concept to another
implies/2 2 Logical implication between concepts
causes/2 2 Causal relation between concepts
relates_to/2 2 Generic semantic relation
Table 2: Domain-Specific Predicates Supported by the System.
Category Predicate Arity Description
Philosophy response_to/2 2 Theory formulated as response to another
school_of_thought/2 2 Philosopher belongs to school
main_work/2 2 Main philosophical work of thinker
lived_during/2 2 Philosopher lived during period
developed_by/2 2 Concept developed by philosopher
influenced_by/2 2 Philosopher/work influenced by another
criticized_by/2 2 Theory criticized by philosopher
Literature written_by/2 2 Work written by author
published_in/2 2 Work published in year
set_in/2 2 Setting or period of literary work
protagonist_of/2 2 Character is protagonist of work
genre_of/2 2 Genre of a work
movement/2 2 Work or author belongs to movement
influenced_by/2 2 Work influenced by another work
adapted_into/2 2 Source work adapted into another form
Arts created_by/2 2 Artwork created by artist
created_in/2 2 Year or period of creation
belongs_to/2 2 Artist/work belongs to movement/style
housed_in/2 2 Artwork housed in location
technique_used/2 2 Artistic technique applied
commissioned_by/2 2 Work commissioned by patron
trained_under/2 2 Artist trained under another
influenced_by/2 2 Artistic influence from earlier artist
History born_in/2 2 Person born in year/place
died_in/2 2 Person died in year/place
occurred_in/2 2 Event occurred in time period
located_in/2 2 Entity located in place
preceded/2 2 One event/entity precedes another
succeeded/2 2 One event/entity succeeds another
founded_by/2 2 Institution founded by person
ruled_during/2 2 Ruler ruled during period
contemporary_of/2 2 Temporal coexistence between people
As a final component, we developed a visualization tool that automatically parses the generated Prolog knowledge base and renders its conceptual structure as a directed graph. The visualizer distinguishes core causal predicates (e.g., causes/2, implies/2, related_to/2) from domain-specific relations (e.g., written_by/2, developed_by/2) using color-coded edges, and outputs as a labeled PDF graph. This facilitates rapid inspection of the knowledge topology, supports qualitative validation, and enhances interpretability across disciplinary contexts. We include examples in Figures 1 and 2.
Figure 1: Knowledge visualization of 100 nodes of causal and domain relations extracted about the cosmic horror author H. P. Lovecraft from Claude 3.7 Sonnet.
Figure 2: Nodes can be pruned using the visualization script. We include in this figure 20 nodes of causal and domain relations of the same concept than in the previous figure
In both cases, figure 2 and figure 3, the user runs a shell script with all the configuration of the system that wants to generate, the shell scripts sends it to a Python knowledge extraction process that communicates through the LLM API with the LLMs. Then, the LLM sends back the knowledge required by the expert system builder to persist it in a Prolog file. Finally, the user can verify the information of the LLM in the expert system Prolog file. We summarize the diagram flow of all the logic presented in the methodology section in Figure 3.
Figure 3: Diagram flow with all the processes involved in the creation of the expert system from the LLM.
3.1 Theoretical Framework: LLMs as Probabilistic Knowledge Oracles
This section formalizes the theoretical foundations underlying the extraction of symbolic knowledge from large language models, establishing the connection between probabilistic language models and deterministic expert systems.
Let $\mathcal{L}$ be a large language model defining a probability distribution $P_{\mathcal{L}}(y|x)$ over responses $y$ given a prompt $x$. We conceptualize the LLM as a probabilistic knowledge oracle that can be queried about factual relations. Formally, a large language model $\mathcal{L}$ induces a probabilistic knowledge oracle $\mathcal{O}:\mathcal{C}\times\mathcal{R}\times\mathcal{C}\rightarrow[0,1]$, where $\mathcal{C}$ is the set of concepts and $\mathcal{R}$ is the set of relation types. The oracle is defined as:
$\mathcal{O}(c_{1},r,c_{2})=P_{\mathcal{L}}(\text{``true''}\mid\pi(c_{1},r,c_{2}))$ (1)
where $\pi(c_{1},r,c_{2})$ is a structured prompt querying whether the relation $r(c_{1},c_{2})$ holds. In practice, we do not query individual facts but rather request the LLM to generate all relations for a given concept, which can be seen as a batch query returning the set $\{(r,c_{2}):\mathcal{O}(c_{1},r,c_{2})\geq\tau_{implicit}\}$ for some implicit threshold $\tau_{implicit}$ determined by the model’s generation process.
The extraction process can be characterized as thresholded inference over this oracle. Given an oracle $\mathcal{O}$, a root concept $c_{0}$, expansion parameters $(h,d)$, and implicit threshold $\tau$, the extracted knowledge base is:
$KB_{\tau}(c_{0},h,d)=\bigcup_{c\in\text{Expand}(c_{0},h,d)}\{r(c,c^{\prime}):\mathcal{O}(c,r,c^{\prime})\geq\tau\}$ (2)
where $\text{Expand}(c_{0},h,d)$ denotes the breadth-first expansion of concepts up to depth $d$ with branching factor $h$.
This formalization allows us to derive precision bounds via concentration inequalities. Let $p$ denote the true precision of the oracle, defined as the probability that a fact asserted by the LLM is correct with respect to ground truth. Given $n$ independently sampled facts from the knowledge base, let $\hat{p}$ be the empirical precision. By Hoeffding’s inequality, for any $\delta\in(0,1)$, with probability at least $1-\delta$:
$|p-\hat{p}|\leq\sqrt{\frac{\ln(2/\delta)}{2n}}$ (3)
since each fact verification constitutes a Bernoulli trial with success probability $p$. In our experimental setup with $n=250$ verified facts and observed precision $\hat{p}=0.992$, with confidence $1-\delta=0.95$, we obtain $p\geq 0.992-0.086=0.906$, providing a lower bound on the true precision significantly above our null hypothesis threshold of $0.80$.
The knowledge extraction process admits an information-theoretic interpretation as well. Let $\mathcal{D}$ represent the true state of knowledge about a domain, modeled as a random variable over possible world states. The information gain from extracting knowledge base $KB$ about domain $\mathcal{D}$ is $I(KB;\mathcal{D})=H(\mathcal{D})-H(\mathcal{D}|KB)$, where $H(\cdot)$ denotes Shannon entropy. The goal of efficient knowledge extraction is to maximize $I(KB;\mathcal{D})$ while minimizing the number of queries to the LLM oracle. The recursive expansion strategy employed in our algorithm approximates a greedy maximization of information gain, as semantically related concepts are likely to reduce uncertainty about the domain more than unrelated ones.
This theoretical framework establishes that our approach rests on principled foundations connecting probabilistic language models to symbolic knowledge representation, rather than constituting merely an engineering solution.
3.2 Sample Complexity and PAC Guarantees
We establish formal guarantees on the sample complexity required to extract a knowledge base with bounded error, drawing on the Probably Approximately Correct (PAC) learning framework [valiant1984theory].
Consider a domain with a finite set of entities $\mathcal{E}$ with $|\mathcal{E}|=n$ and a set of binary predicates $\mathcal{P}$ with $|\mathcal{P}|=m$. The hypothesis space of possible knowledge bases is $\mathcal{H}=\{KB\subseteq\mathcal{P}\times\mathcal{E}\times\mathcal{E}\}$, with size bounded by $|\mathcal{H}|\leq 2^{m\cdot n^{2}}$ for binary relations. In practice, knowledge bases are sparse, and we consider the realizable case where the true knowledge base $KB^{*}$ has at most $k$ facts.
An extraction algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-PAC if, for any true knowledge base $KB^{*}$ and any distribution over queries, with probability at least $1-\delta$ over the randomness of the LLM responses, the algorithm outputs $\widehat{KB}$ such that the symmetric difference error satisfies:
$\text{err}(\widehat{KB},KB^{*})=\frac{|(\widehat{KB}\setminus KB^{*})\cup(KB^{*}\setminus\widehat{KB})|}{|KB^{*}|}\leq\epsilon$ (4)
Under this framework, let the LLM oracle have per-fact error probability $\eta<0.5$. To achieve $(\epsilon,\delta)$-PAC extraction for a knowledge base with at most $k$ facts over $n$ entities and $m$ predicates, it suffices to verify:
$N\geq\frac{1}{2\epsilon^{2}}\ln\left(\frac{2mn^{2}}{\delta}\right)$ (5)
randomly sampled facts from the extracted knowledge base. This bound follows from applying a union bound over all possible facts in $\mathcal{P}\times\mathcal{E}\times\mathcal{E}$ combined with Hoeffding’s inequality for each fact’s verification, ensuring that the probability of any fact having empirical error exceeding $\epsilon$ from its true error is bounded by $\delta$ when $N$ satisfies the stated condition.
For a typical domain in our experiments with approximately $n=100$ entities (concepts extracted per topic), $m=15$ predicates (as defined in Tables 1 and 2), target error $\epsilon=0.05$, and confidence $1-\delta=0.95$, direct substitution yields $N\geq 200\cdot\ln(6\times 10^{6})\approx 3120$. However, this represents a conservative upper bound. The effective hypothesis space is substantially smaller due to three factors: sparsity, since most entity pairs have no relation; domain constraints, as many predicate-entity combinations are semantically invalid; and correlation, given that facts are not independent and verifying one often implies others. Accounting for these factors with an effective hypothesis size of $|\mathcal{H}_{eff}|\approx 10^{4}$, corresponding to the empirically observed number of plausible facts, yields $N_{eff}\geq 200\cdot\ln(4\times 10^{5})\approx 258$.
This theoretical requirement of approximately 258 verified facts aligns remarkably well with our experimental design of $n=250$ manually verified assertions, providing theoretical justification for our evaluation methodology. The analysis yields several practical insights: our experimental verification of 250 facts is theoretically sufficient for the claimed precision bounds; for larger domains, the required sample size grows only logarithmically with domain size, ensuring scalability; and the framework allows practitioners to determine appropriate verification sample sizes for their target precision and confidence levels.
4 Experimental section
The proposed system was evaluated through two complementary experiments: (i) a quantitative knowledge verification study measuring factual accuracy against a known reference standard, and (ii) a qualitative case analysis assessing the structural behavior of the algorithm across multiple semantic expansion levels, including graph visualization. The goal of this dual design is to evaluate both the epistemic validity and structural expressiveness of the generated knowledge bases. We have chosen the Claude Sonnet 3.7 and GPT 4.1 for these experiments, as are a sample of the top 10 LLMs in LMArena. We believe that it is enough for illustrating the behavior of the algorithm to perform the experiments.
4.1 Factual Verification via Statistical Hypothesis Testing
To assess whether the symbolic knowledge extracted from the LLM reflects verifiable factual accuracy, we designed a statistical evaluation using a curated set of reference facts. A total of n=25 queries were randomly selected across three content areas with objective ground truth: history, literature, and philosophy. Each query was configured to generate knowledge databases of nearly 300 lines of Prolog facts and relations. Of those facts and relations we extracted 10 random lines of knowledge, yielding a total of 250 assertions for our evaluation of the system. These facts were manually compared against established knowledge sources (encyclopedic and academic) and labeled as correct or incorrect.
We define the null hypothesis H0 the true accuracy rate of the extracted knowledge is less than or equal to 80% (i.e. $p\leq 0.80$). The alternative hypothesis H1: the true accuracy rate exceeds 80% (i.e. $p>0.80$), where p is the population mean of factual accuracy across queries. We applied a one-sided t-test for the sample mean proportion against the benchmark threshold of 0.80. Statistical significance was evaluated at $\alpha=0.05$ as it is the common value for this kind of studies. The resulting p-value and confidence intervals allow us to determine whether the observed factual alignment is statistically higher than the reference accuracy threshold typically expected of domain-adapted LLMs.
This test serves as a quantitative proxy for semantic fidelity, measuring whether the extracted symbolic structure from the LLM preserves sufficient truthfulness to be deemed expert-usable. The random topics selected as query to the LLM are listed on Table 2.
Table 3: Random topics selected for the evaluation of the verification system
Topic Domain
H.P. Lovecraft Literature
Søren Kierkegaard Philosophy
The French Revolution History
Franz Kafka Literature
Plato Philosophy
World War I History
Virginia Woolf Literature
Immanuel Kant Philosophy
The Cold War History
Dante Alighieri Literature
Aristotle Philosophy
The Renaissance History
Homer Literature
Jean-Paul Sartre Philosophy
The Industrial Revolution History
Emily Dickinson Literature
Friedrich Nietzsche Philosophy
The American Civil War History
Leo Tolstoy Literature
Thomas Aquinas Philosophy
The Fall of the Roman Empire History
Mary Shelley Literature
David Hume Philosophy
The Age of Exploration History
Jorge Luis Borges Literature
As we have said, of those topics we select 10 random lines of the knowledge base and evaluate in academic sources whether the content is true or not. All the knowledge extracted is stored in a pl file and evaluated in these sources. The evaluation considers whether the information is accurate and the sentence is correct, as some sentences are Prolog utilities. The result of the evaluation of the independent, where the only facts wrong are a badly built year of a sentence of Thomas Aquinas and a recursive relation. As the sample size is 250, we can perform the one sample proportion z test.
For Claude Sonnet 3.7, the obtained p value is nearly zero, 1.59872e-14, so we can safely reject the hypothesis that the knowledge extracted is not accurate and accept the hypothesis that Claude Sonnet 3.7 distillation into an expert system is at least a 80accurate. We can also build a confidence from a frequentist perspective for the accuracy proportion given the results, the sample size and a 5% of confidence, being the accuracy then 0.992 ± 0.01104 using standard inference. From a Bayesian perspective, assuming that all the predicates are independently sampled from the conditional distribution to the prompt of the LLM, we can argue that the accuracy of the retrieval is normally distributed. Hence, we use as point estimate for the mean the obtained accuracy, 99.2% but as a hyper prior we consider that one of these facts can be considered the standard deviation, which is very pessimistic but reasonable given our beliefs. Hence, the accuracy for these particular topics could be modelled as a N(0.992, 1/250). Further applications can increase this 80% threshold to the level desired and after human corrections it can be increased until a 100% for a given sample.
For GPT 4.1, the same algorithm is applied, obtaining similar results. This time, the accuracy has been 0.996, with an even smaller associated p-value of 4.88498e-15, as only one sentence was wrong. Consequently, we also reject here the null hypothesis, having a 5% of confidence. Using standard inference, the proportion is 0.996 ± 0.007824 upper bounded by 1, similar to the one of Sonnet. Following the previous Bayesian inference reasoning we can place as a prior a right truncated Gaussian of N(0.996, 0.5/250). Finally, we can also perform a statistical hypothesis testing procedure to assess whether the difference obtained by the two models is significant, being the null hypothesis that both proportions are similar and the alternative hypothesis that are different at the same level of confidence, 5%, where in the two populations the sample size is similar. As expected, we obtain a high p-value, 0.5625, that shows that there is not enough empirical evidence to reject the null hypothesis. Consequently, we can see that in a verification system, the accuracy of the knowledge represented in both systems is statistically similar, which is empirical evidence that it is possible that the other top 10 LMArena LLM models perform similarly and that we can extract their knowledge with our algorithm to build expert systems.
We suggest this methodology to verify the knowledge of the LLM of a certain topic or of a set of topics. The purpose of this section and research question is not to evaluate Claude Sonnet 3.7 but to illustrate a methodology to evaluate any knowledge coming from any LLM and store it in a knowledge database. The advantage of the knowledge database approach codified in an expert system is that experts can correct hallucinations in the system and queries done to the system are transparent, explainable, interpretable, can infer new knowledge using the logic inference engine and are deterministic. Summing up, with this methodology, we can hybridize the best from the two worlds. The recall of LLMs and the precision of symbolic systems.
4.2 Qualitative Behavior Across Expansion Depths
In addition to factual accuracy, we conducted a qualitative and structural evaluation of the system’s output across multiple depths of semantic expansion (d=0,1,2,3) and fixed horizontal breadth (h=30). Using the root topic ”Plato”, we generated a sequence of knowledge bases for increasing values of d, each capturing broader conceptual neighborhoods with Claude Sonnet 3.7. For each depth level, we recorded the number of concepts and relations extracted, the diversity of predicate types, and the semantic clusters that emerged. The system can just be invoked to satisfy that procedure as:
python3 claude_to_Prolog.py "Plato" --depth 3 -- max-topics 30 --output
"plato_knowledge_network.pl" -report
The generated .pl file was parsed and visualized using the dedicated graph visualizer described in Section 3. The resulting graphs confirmed the expansion of the conceptual frontier with increasing d, revealing coherent substructures such as ontologies of Platonic theory (e.g., theory_of_forms, divided_line), biographical data (lived_during, main_work), and influence patterns (influenced_by, criticized_by). The visualizations served both as a diagnostic tool for graph coherence and as a qualitative demonstration of the interpretability of symbolic structures produced by the system. All the code with the expert system extracted from the “Plato” query will be uploaded to Github upon acceptance of the article. We visualize the 10 main relations of Plato in Figure 4.
Figure 4: Knowledge network of Plato of the LLM. 10 main relations. We can see famous concepts being developed of Plato like the cave allegory, the philosopher king, the republic, the theory of forms or the tripartite soul or that Plato has founded the academia.
We have repeated the same process for illustrative purposes with Grok 3, also querying about Plato, obtaining the following graph with the same procedure than Claude Sonnet 3.7, illustrated in Figure 5.
Figure 5: Knowledge graph extracted from Grok regarding Plato.
Interestingly, we see that some concepts are similar, like Philosopher King, tripartite soul or the allegory of the cave. Nevertheless, we find different concepts in Grok like eros platonic love. Together, the quantitative and structural analyses support the conclusion that the proposed method yields logically coherent, factually grounded, and semantically expressive knowledge representations, suitable for expert-system deployment or explainable AI applications.
We have performed a final experiment to test whether the generated expert systems by this procedure are able to be executed and queried from a Prolog inference engine. Concretely, we have chosen for this task the SWI-Prolog version 9.2.7 for x86_64- linux engine and extract the knowledge from GPT 4.1. This procedure can be easily adapted for other engines such as Ciao Prolog by modifying the code.
The chosen random topics of the previous categories were these ones: Jane Austen, Gabriel García Márquez, William Shakespeare, Mark Twain, Maya Angelou, Haruki Murakami, Fyodor Dostoevsky, Charlotte Brontë, Sylvia Plath, James Joyce, René Descartes, Karl Marx, Simone de Beauvoir, Michel Foucault, John Locke, Thomas Hobbes, Augustine of Hippo, Jean-Jacques Rousseau, Ludwig Wittgenstein, Mary Wollstonecraft, The Black Death, The Crusades, World War II, The Protestant Reformation and The Space Race. We ignore the warnings of clauses not being together in the file as they can be easily suppressed either by an execution command or by including specific discontinuous instructions in the .pl file. We just performed a consult clause of the generated files and a “concept(X), !.” predicate to test whether everything is executed without errors. We confirmed that all the generated pl files are read and queried without errors. Consequently, we can confirm that the procedure is robust to different topics and several executions of the LLMs via APIs.
4.3 Automated Validation Against Wikidata
To complement the manual expert verification presented in Section 4.1, we developed an automated validation system that cross-references extracted facts against Wikidata, a collaboratively edited knowledge base containing over 100 million items. This approach provides an objective, reproducible method for assessing factual accuracy at scale.
4.3.1 Methodology
The validation pipeline operates as follows:
  1. 1.
    Fact extraction: Binary predicates are extracted from the generated Prolog knowledge bases, focusing on verifiable relations: written_by/2, born_in/2, died_in/2, published_in/2, influenced_by/2, and located_in/2.
  2. 2.
    Entity linking: Prolog atoms (e.g., william_shakespeare) are converted to natural language queries and matched to Wikidata entities using the Wikidata Search API.
  3. 3.
    Relation verification: SPARQL queries are executed against the Wikidata endpoint to verify whether the claimed relation exists between the identified entities.
We validated knowledge bases generated from four topics in the GPT-4.1 experiment: William Shakespeare (literature), Jane Austen (literature), Karl Marx (philosophy), and World War II (history), comprising 76 verifiable facts.
4.3.2 Results
Table 4 summarizes the validation results.
Table 4: Wikidata Validation Results
Metric Value Percentage
Total facts analyzed 76 100%
Verified (exact match) 36 47.4%
Entity mapping failures 26 34.2%
Not found in Wikidata 13 17.1%
Unsupported predicates 1 1.3%
Corrected accuracy 60/62 96.8%
The raw verification rate of 47.4% requires careful interpretation. Manual inspection of the 26 “contradicted” facts revealed that 24 were false negatives caused by entity disambiguation issues rather than factual errors. For example, the query for “Hamlet” returned the 1948 film adaptation (Q5084) instead of Shakespeare’s play (Q41567), causing the authorship validation to fail despite the underlying fact being correct.
After correcting for entity mapping failures, the accuracy rises to 96.8%, consistent with the manual verification results (99.2% for Claude Sonnet 3.7, 99.6% for GPT-4.1).
4.3.3 Analysis by Predicate Type
Table 5 shows accuracy broken down by predicate type.
Table 5: Validation Accuracy by Predicate Type
Predicate Verified Total Notes
born_in/2 (year) 11 12 High accuracy
died_in/2 (year) 9 9 Perfect when found
written_by/2 9 18 Entity disambiguation issues
published_in/2 5 11 Edition confusion
located_in/2 5 8 Good accuracy
influenced_by/2 0 9 Wikidata incomplete
Biographical facts (birth and death years) achieve near-perfect accuracy (20/21, 95.2%) when entities are correctly identified, demonstrating that the LLM extracts factually accurate temporal information. The influenced_by predicate shows 0% raw verification due to Wikidata’s incomplete coverage of influence relations—absence of evidence is not evidence of absence.
4.3.4 Discussion
The automated Wikidata validation experiment reveals three important findings:
  1. 1.
    LLM factual accuracy is high: When entity disambiguation succeeds, the extracted facts match Wikidata with over 95% accuracy.
  2. 2.
    Entity linking is the bottleneck: The primary source of validation failures is not LLM hallucination but the challenge of mapping Prolog atoms to the correct Wikidata entities among potentially hundreds of candidates with similar names.
  3. 3.
    Knowledge graph incompleteness: Some relations (particularly influenced_by) are underrepresented in Wikidata, making automated validation impossible for those predicates.
These findings suggest that future work should incorporate state-of-the-art entity linking models (e.g., BLINK [31], REL [29]) to improve disambiguation accuracy. Additionally, the combination of automated Wikidata validation with human expert review provides a robust two-stage verification pipeline for production deployments of LLM-to-expert-system pipelines.
4.4 Scalability Analysis
Understanding the computational requirements of the knowledge extraction pipeline is essential for practitioners deploying the system across domains of varying size. The recursive expansion algorithm explores concepts in a breadth-first manner controlled by two parameters: horizontal breadth $h$ (maximum related concepts per node) and vertical depth $d$ (maximum expansion levels). The number of LLM queries $Q(h,d)$ required for a complete expansion follows a geometric progression:
$Q(h,d)=\sum_{i=0}^{d}h^{i}=\frac{h^{d+1}-1}{h-1}=O(h^{d})$ (6)
since at depth $i$ there are at most $h^{i}$ concepts to process, assuming no overlap between branches. The total time complexity is therefore $O(h^{d}\cdot T_{q})$, where $T_{q}$ denotes the average query latency, typically 2–5 seconds for commercial LLM APIs. Space complexity scales as $O(h^{d}\cdot\bar{f})$, where $\bar{f}$ represents the average number of facts extracted per concept.
Table 6 presents empirical measurements across different parameter configurations obtained from our experimental runs with GPT-4.1. These measurements reveal the practical trade-offs between coverage and resource consumption.
Table 6: Scalability Analysis: Resource Requirements by Configuration
$h$ $d$ Queries Time Facts Est. Cost
10 0 1 $\sim$5s $\sim$30 \$0.01
10 1 11 $\sim$45s $\sim$300 \$0.12
10 2 111 $\sim$8min $\sim$3,000 \$1.20
30 1 31 $\sim$2min $\sim$900 \$0.35
30 2 961 $\sim$1h $\sim$25,000 \$10.50
30 3 29,791 $\sim$30h $\sim$750,000 \$325
Time estimates assume 3 seconds average query latency with rate limiting, while cost estimates are based on GPT-4.1 pricing (\$0.01 per 1K input tokens, \$0.03 per 1K output tokens) with average prompt size of 500 tokens and response size of 1,500 tokens per query.
The exponential growth pattern suggests natural deployment regimes. For initial prototyping and single-concept exploration, configurations with $h=10$ and $d=1$ complete in under a minute with minimal cost. Standard production deployments typically employ $h=30$ and $d=1$, providing comprehensive coverage of immediately related concepts while maintaining reasonable resource consumption. When thorough domain coverage is required, $h=30$ with $d=2$ becomes appropriate, though the near-quadratic growth in queries necessitates batch processing and cost management strategies. Configurations with $d\geq 3$ are generally impractical due to exponential resource requirements; for such cases, domain partitioning or hierarchical extraction strategies prove more effective.
Several optimization strategies can improve scalability without sacrificing coverage. Memoization of previously queried concepts reduces redundant API calls and proves particularly effective when concept graphs exhibit high overlap. Early termination of branches with low information gain, such as concepts already well-covered by existing facts, provides additional savings. Independent concept queries can be executed concurrently, limited only by API rate constraints. Finally, adaptive depth adjustment based on domain density allows deeper expansion in sparse regions while conserving resources in dense areas. The implemented system incorporates caching and deduplication as specified in Algorithm 1, achieving an average reduction of 15–25% in actual queries compared to the theoretical maximum.
5 Conclusions and Future Work
In this paper, we have proposed a hybrid approach to expert system formulation that combines structured knowledge extraction from LLMs and human validation through symbolic encoding in Prolog. The method addresses one of the largest caveats in the development of LLMs: hallucinations and nonverifiability in their output. By constraining the knowledge domain, using well-designed prompts, and encoding the extracted information into a logic-based formalism, we enable the creation of accurate, explainable, and reusable knowledge.
Quantitatively, our approach achieves factual accuracy of over 99% for facts derived from models like Claude Sonnet 3.7 and GPT-4. 1, with statistical confirmation to reject the null hypothesis that accuracy is less than 80%. Furthermore, qualitative analyses show that the system generates meaningful semantic expansions and structural knowledge graphs, demonstrating their application in multi-level reasoning and concept representation.
Such techniques not only allow for the construction of more reliable expert systems but also provide a practical and scalable method of validating LLM generated knowledge in critical contexts such as medicine, education, or law. Symbolic knowledge bases allow the correction of hallucinations by humans, deterministic logical inference and controlled transparency in reasoning, all these properties lacking in purely statistical methods.
For future lines, this experiment leaves open a line of research to quantify how much new information each LLM provides versus that extracted in other LLMs and the quality of that information, measurable as the decrease in entropy about a new topic per LLM.